# TimeAutoDiff: Combining Autoencoder and Diffusion model for time series tabular data synthesizing

### v1

@@ 4.1 Metrics for Evaluations
For quantitative evaluation of synthesized data, we mainly focus on three criteria (1) distributional similarities of two tables; (2) the usefulness for predictive purposes; (3) the temporal and feature dependencies; We employ the following evaluation metrics:
@@
    Predictive Score [39] measures the utility of generated sequences by training a posthoc sequence prediction model (optimizing a 2-layer LSTM) to predict next-step temporal vectors under Train-on-Synthetic-Test-on-Real (TSTR) framework.
@@
    Temporal Discriminative Score measures the similarity of distributions of inter-row differences between generated and original sequential data. This metric is designed to see if the generated data preserves the temporal dependences of the original data. For any fixed integer $t\in\{1,\dots,T-1\}$, the difference of $n$-th row and $(n+t)$-th row in the table over $n\in\{1,\dots,T-t\}$ is computed for both generated and original data, and discriminative score [39] is computed over the differenced matrices from original and synthetic data. We average discriminative scores over $10$ randomly selected $t\in\{1,\dots,T-1\}$.
@@ 4.2 Experimental Results
For single-sequence datasets, it significantly improves the (temporal) discriminative and feature correlation scores in all $6$ datasets over the baseline models. We train the RNN to predict a certain column in the dataset to measure the predictive score. The columns predicted in each dataset are listed in Table $4$ in the Appendix B. TimeAutoDiff also dominates the predictive score metric. But for some datasets, the performance gaps with the second-best model are negligible, for instance, TSGM for Hurricane and AirQuality datasets. It is intriguing to note that predictive scores can be good even when we have low-fidelity guarantees. For multi-sequence datasets, (C-)TimeAutoDiff shows the best results in almost all cases under all $4$ metrics, except for the predictive score in the nasdaq100 dataset. Nonetheless, the performance differences between (C-)TimeAutoDiff and the best model are minimal for both metrics. We did not observe significant differences between conditional and unconditional generations in the datasets we considered.
@@
Summary of Table 3: TimeAutoDiff is characterized by multiple components in both the VAE and diffusion models. The components include: (1) sinusoidal encodings of numerical features; i.e., Eq. (1) (VAE), (2) transformer encoder block with two 1D convolution layers (VAE), (3) the encodings of timestamp information (diffusion model), and (4) the Bi-directional RNN (diffusion model). Experimental results are summarized in Table 3 with the model that shows the best performances bolded. No model dominates four metrics over the four single-sequence datasets. Nonetheless, TimeAutoDiff demonstrates satisfactory performance in 7 out of 16 cases. Even in the majority of the remaining 9 cases, the performance gaps compared to the best model are negligible. This observation implies all the $4$ components are necessary. The inclusion of components related to the diffusion model, such as timestamp encoding and Bi-directional RNN, significantly impacts the generative performance across all cases as models lacking these components do not exhibit optimal performance. The inclusion of encodings for numerical features in the VAE notably enhances the fidelity and temporal dependencies of the generated data.
@@ 5 Future Directions
TimeAutoDiff focuses on ensuring the good fidelity and utility of the synthesized data. However, the privacy guarantees of the model is not extensively explored in the current manuscript, despite its potential critical importance in domains such as healthcare and finance. With recent efforts to ensure differential privacy guarantees on latent diffusion model-based tabular data synthesizers [46], we suggest a similar approach could be applied to TimeAutoDiff.

### v2

@@ 4.1 Metrics for Evaluations
For the quantitative evaluation of synthesized data, we mainly focus on three criteria (1) the distributional similarities of the two tables; (2) the usefulness for predictive purposes; (3) the temporal and feature dependencies; We employ the following evaluation metrics:
@@
    Predictive Score [39] measures the utility of generated sequences by training a posthoc sequence prediction model (optimizing a 2-layer LSTM) to predict next-step temporal vectors under a Train-on-Synthetic-Test-on-Real (TSTR) framework.
@@
    Temporal Discriminative Score measures the similarity of distributions of inter-row differences between generated and original sequential data. This metric is designed to see if the generated data preserves the temporal dependencies of the original data. For any fixed integer $t\in\{1,\dots,T-1\}$, the difference of the $n$-th row and $(n+t)$-th row in the table over $n\in\{1,\dots,T-t\}$ is computed for both generated and original data and discriminative score [39] is computed over the differenced matrices from original and synthetic data. We average discriminative scores over $10$ randomly selected $t\in\{1,\dots,T-1\}$.
@@ 4.2 Experimental Results
For single-sequence datasets, it significantly improves the (temporal) discriminative and feature correlation scores in all $6$ datasets over the baseline models. We train the RNN to predict a certain column in the dataset to measure the predictive score. The columns predicted in each dataset are listed in Table $4$ in Appendix B. TimeAutoDiff also dominates the predictive score metric. But for some datasets, the performance gaps with the second-best model are negligible, for instance, TSGM for Hurricane and AirQuality datasets. It is intriguing to note that predictive scores can be good even when we have low-fidelity guarantees. For multi-sequence datasets, (C-)TimeAutoDiff shows the best results in almost all cases under all $4$ metrics, except for the predictive score in the nasdaq100 dataset. Nonetheless, the performance differences between (C-)TimeAutoDiff and the best model are minimal for both metrics. We did not observe significant differences between conditional and unconditional generations in the datasets we considered.
@@
Summary of Table 3: TimeAutoDiff is characterized by multiple components in both the VAE and diffusion models. The components include: (1) sinusoidal encodings of numerical features; i.e., Eq. (1) (VAE), (2) transformer encoder block with two 1D convolution layers (VAE), (3) the encodings of timestamp information (diffusion model), and (4) the Bi-directional RNN (diffusion model). Experimental results are summarized in Table 3 with the model that shows the best performances in bolded. No model dominates four metrics over the four single-sequence datasets. Nonetheless, TimeAutoDiff demonstrates satisfactory performance in 7 out of 16 cases. Even in the majority of the remaining 9 cases, the performance gaps compared to the best model are negligible. This observation implies all the $4$ components are necessary. The components related to the diffusion model, such as timestamp encoding and Bi-directional RNN, significantly impact the generative performance across all cases as models lacking these components do not exhibit optimal performance. The encodings for numerical features in the VAE notably enhance the fidelity and temporal dependencies of the generated data.
@@ 5 Future Directions
TimeAutoDiff focuses on ensuring the good fidelity and utility of the synthesized data. However, the privacy guarantees of the model are not extensively explored in the current manuscript, despite its potential critical importance in domains such as healthcare and finance. With recent efforts to ensure differential privacy guarantees on latent diffusion model-based tabular data synthesizers [46], we suggest a similar approach could be applied to TimeAutoDiff.